OpenStudy (dls):

If (x+iy)^3=u+iv then show that..............................

OpenStudy (dls):

\[\LARGE \frac{u}{x}+\frac{v}{y}=4(x^2-y^2)\]

OpenStudy (dls):

@shubhamsrg @ParthKohli @Callisto

OpenStudy (shubhamsrg):

expand LHS and compare maybe ?

OpenStudy (dls):

\[\LARGE x^3-y^3-3xy^2+3x^2iy=u+iv\]

OpenStudy (dls):

3x^iy=iv?

OpenStudy (dls):

3x^2iy=iv*?

Parth (parthkohli):

Just a thought: \(N(x + yi) = x^2 + y^2\). So \(N((x + yi)^3) = (x^2 + y^2)^3\). Also, \(N (u + vi) = u^2 + v^2\)\[(x^2 + y^2)(x^2 + y^2)(x^2 + y^2) = u^2 + v^2\]

Parth (parthkohli):

I don't know if that's the method heh.

OpenStudy (dls):

got it guys thanks,no not you @ParthKohli :P

Parth (parthkohli):

lol

OpenStudy (anonymous):

lol

OpenStudy (dls):

\[\LARGE x^3-iy^3+3x^2iy+3x i^2y^2\]

Parth (parthkohli):

Sahi hai, last two terms factor kar lo.

Parth (parthkohli):

Ohhh, got it! @shubhamsrg was right... compare :-P

OpenStudy (shubhamsrg):

someone call @yrelhan4 :/

OpenStudy (dls):

yeah compared and gotcha

OpenStudy (dls):

u=x^3-3xy^2

Parth (parthkohli):

(y)

OpenStudy (dls):

why so many people???please

OpenStudy (dls):

embarrasing

Parth (parthkohli):

These questions are interesting. :-)

OpenStudy (anonymous):

nahhh don't mind us, go on

OpenStudy (shubhamsrg):

log(x+iy) , in case you'd want to try that way ?

OpenStudy (dls):

WHAT LOG WHY

OpenStudy (shubhamsrg):

just an alternative method, hmm, you may ignore it though, nvm! ^_^

OpenStudy (dls):

ignored

OpenStudy (anonymous):

yrelhan4 ka time khatam. its time to RnR :P

OpenStudy (dls):

yuhuuuuuuuuuuuuu :P

OpenStudy (anonymous):

Aur meri itni besti mat kiya kar.. :/ @shubhamsrg

OpenStudy (dls):

rock n roll :P

OpenStudy (shubhamsrg):

ooo teri! new id! :O

OpenStudy (dls):

we have a winner her e<3

OpenStudy (anonymous):

haan.. medal vaapis le shivam :/

OpenStudy (dls):

no :P

OpenStudy (dls):

lol 2 medals

OpenStudy (anonymous):

Hmm. Thank you! :/

OpenStudy (dls):

http://i.qkme.me/35btfs.jpg

OpenStudy (shubhamsrg):

LAWL ! :D

OpenStudy (anonymous):

Hmm. :/

OpenStudy (anonymous):

Chalo bye.

OpenStudy (dls):

:)